ON A SUBFACTOR CONSTRUCTION OF A FACTOR NON ANTI-ISOMORPHIC TO ITSELF

MARIA GRAZIA VIOLA

Abstract. We give a subfactor construction for a II_{1} factor M which is not antiisomorphic to itself. The II_{1} factor we consider is essentially the same as the example previously given by Connes. However, our construction uses the recently developed theory of free group factors. We show that there exists an inclusion of II_{1} factors A ⊂ B which by iteration of the Jones basic construction produces M as the enveloping algebra. Here A is a free group factor and B is isomorphic to the crossed product of A by an action of a finite group. By using a Connes’ argument involving the invariant χ(M), we verify that M is not anti–isomorphic to itself.